Question 4

Easy Decorating uses a job-order costing system to collect the costs of its interior decorating business. Each client's consultation is treated as a separate job. Overhead is applied to each job based on the number of decorator hours incurred. List below are data for the current year:

Estimated overhead $840,000

Actual overhead $861,000

Estimated decorator hours 40,000

Actual decorator hours 41,500

Instructions

(a) Compute the predetermined overhead rate: $______ per decorator hour.

(b) Determine whether the overhead was under or over applied and by how much.
Answer #1

a) calculation of predetermined overhead rate:

Predetermined overhead rate= estimated overhead/estimated decorator hours

Predetermined overhead rate= 840000/40000= 21

Predetermined overhead rate= 21

b) actual overhead rate= 861000/41500= 20.75

Overhead is overapplied= 21-20.75= 0.25